# The Progenitor Systems of Classical Novae in M31

### v1

@@ 2 Data
Our work builds off decades of observations of M31; namely, a historical nova catalog covering the entire galaxy, originally created for comparison with an X-ray nova catalog (Pietsch et al., 2007), but regularly updated in a publicly available website maintained by W. Pietsch111https://www.mpe.mpg.de/m31novae/opt/m31/. We combine these data with a spatially resolved ancient star formation history of M31 (Williams et al., 2017) produced from the PHAT survey (Dalcanton et al., 2012; Williams et al., 2014). This map provides the crucial link between the locations of novae and the age distribution of the stars surrounding them, enabling the recovery of a DTD.
@@ 2.2 PHAT stellar age distribution map
The PHAT survey collected HST photometry for 117 million individual stars in the northern half of M31, measuring the stars simultaneously in 6 bands with coverage from the UV to the IR. This rich data set, hosted on MAST at http://dx.doi.org/10.17909/T91S30 (catalog 10.17909/T91S30), has resulted in dozens of publications that focus on different aspects of the stellar content of M31. Williams et al. (2017) recovered the star formation history by dividing the PHAT data into 826 (deprojected to 0.31.4 kpc at the distance of M31) spatial cells, producing color-magnitude diagrams (CMDs) for each cell that were then fit with the MATCH software package (Dolphin, 2002, 2012, 2013), taking into account the spatially varying extinction inferred from a high resolution map by Dalcanton et al. (2015).
Each cell was fit using four different sets of isochrones: Padova (Marigo et al., 2008; Girardi et al., 2010), BaSTI (Pietrinferni et al., 2004; Cassisi et al., 2005; Pietrinferni, Adriano et al., 2013), PARSEC (Bressan et al., 2012), and MIST (Choi et al., 2016). This resulted in four maps of the SAD (or star formation history) in the PHAT footprint, one for each set of isochrones, each with 16 temporal bins spanning from the present day to the age of the Universe (see Williams et al. 2017 for details). Williams et al. accounted for chemical enrichment by assuming that the metal content in M31 has a mean that decreases exponentially with lookback time with a constant spread. They fit the time scale of the exponential separately for the outer disk, inner disk, and bulge of M31.
@@ 5 Discussion and Conclusions
Taken together, our detections and upper limits are consistent with a decaying rate of nova formation with delay time; though, strictly speaking, we cannot rule out a uniform formation rate. A decaying formation rate is borne out by simple arguments in stellar evolution. Technically, a binary system can produce novae as soon as the more massive partner evolves into a white dwarf. As delay time increases, the zero-age main sequence masses of both the donor and the degenerate accretor decrease on average, leading to a decrease in formation efficiency of novae (Ritter et al., 1991; Kolb, 1995; Chen et al., 2016) and an increase in recurrence time (Yaron et al., 2005), as less mass is available to transfer to the accretor. Although the PHAT SAD map lacks sufficient resolution at short delay times to probe the “turn-on” phase – the prediction that the formation efficiency of novae should be zero before the formation of the first WDs at delays of 40 Myr –, it does probe, and successfully recover, the broad expectation of a decrease in formation efficiency at longer delays.
Our rough completeness corrections allow us to estimate nova rates (per unit stellar mass) as well as nova counts; these rates are consistent with the shape and height of the Kemp et al. DTD (see Figure 6), with the effective survey length correction yielding a slightly better match than the time-limited sample. To be clear, no adjustments have been made to our data to bring them into alignment with the theoretical rates. This agreement is simply a product of the DTD recovery process outlined in Section 3 and the completeness corrections presented in Section 2.1.
@@
We conclude with a brief discussion of the comparison between our nova DTD and the DTD for Type Ia SNe recovered by Maoz et al. (2012), which is interesting as some subclasses of novae have been proposed as potential SN Ia progenitors (Maoz et al., 2014). Our completeness-corrected nova DTD is about four orders of magnitude higher than the SN Ia DTD, and noticeably shallower at long delay times. This implies that if a subclass of nova-producing systems do go on to explode as SN Ia, they must be a very small fraction () of the classical nova systems that explode in star-forming galaxies like M31.
Our work reinforces the previously demonstrated power of DTDs for constraining the progenitor populations of the products of binary stellar evolution. The continuation of high-cadence surveys of nearby, well-studied galaxies such as M31 would provide better constraints on the nova DTD at all lookback times and reduce our reliance on historical nova catalogs of dubious completeness. Resolved stellar populations of those galaxies provide opportunities for the recovery of DTDs for other astrophysical transients. SAD maps similar to the one we used for M31 exist for other galaxies in the Local Group (e.g., Lazzarini et al. 2022 for M33, Mazzi et al. 2021; Harris & Zaritsky 2009 for the LMC, Rubele et al. 2018; Harris & Zaritsky 2004 for the SMC), with matching, high-quality catalogs of astrophysical transients and other products of stellar evolution. This research enables further application of the methods described here, which could be leveraged to constrain BPS models, deepening our understanding of key stages of binary stellar evolution and improving the predictions for the rates of rare events such as SNe Ia and black hole mergers.
@@
We acknowledge fruitful discussions with Dan Maoz, Peter Nugent, and Robin Ciardullo. We are grateful to Alex Kemp, who kindly provided us with a digital form of his nova DTD. L. Galbany acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, and 2021-SGR-01270. This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided. Specifically, this work used the HTC cluster, which is supported by NIH award number S10OD028483. We acknowledge financial support from NSF grant AST-2307865, and HST grants HST-GO-16741.002-A, HST-GO-16287.002-A, and HST-GO-17179.002-A.

### v2

@@ 2 Data
Our work builds off decades of observations of M31; namely, a historical nova catalog covering the entire galaxy, originally created for comparison with an X-ray nova catalog (Pietsch et al., 2007), but regularly updated in a publicly available website maintained by W. Pietsch111https://www.mpe.mpg.de/~m31novae/opt/m31/. We combine these data with a spatially resolved ancient star formation history of M31 (Williams et al., 2017) produced from the PHAT survey (Dalcanton et al., 2012; Williams et al., 2014). This map provides the crucial link between the locations of novae and the age distribution of the stars surrounding them, enabling the recovery of a DTD.
@@ 2.2 PHAT stellar age distribution map
The PHAT survey collected HST photometry for 117 million individual stars in the northern half of M31, measuring the stars simultaneously in 6 bands with coverage from the UV to the IR. This rich data set, hosted on MAST at http://dx.doi.org/10.17909/T91S30 (catalog 10.17909/T91S30) (Dalcanton, Julianne & Williams, Benjamin, 2012), has resulted in dozens of publications that focus on different aspects of the stellar content of M31. Williams et al. (2017) recovered the star formation history by dividing the PHAT data into 826 (deprojected to 0.31.4 kpc at the distance of M31) spatial cells, producing color-magnitude diagrams (CMDs) for each cell – assuming a Kroupa initial mass function (Kroupa, 2002) – that were then fit with the well-tested MATCH software package (Dolphin, 2002). MATCH also contains tools to model both systematic (Dolphin, 2012) and random (Dolphin, 2013) uncertainties, which the authors implement when deriving their SAD map.
Certain features in a CMD are clear indicators of the presence of stars of a certain age; by comparing an observed CMD to a model composed of simple stellar populations of various ages, an SAD can be robustly recovered. Williams et al. (2017) used MATCH 2.6 to calculate the best-fit SADs to CMDs drawn from several bands, only including stars brighter than the 50% completeness limit and making use of the software’s ability to incorporate complex extinction models. They draw from a high-resolution dust map of M31 (Dalcanton et al., 2015) derived from a two-component model of reddening-induced spread among RGB stars. This model provides a galactic map of median , spread of the log-normal distribution, and fraction of stars being affected by dust; Williams et al. (2017) then feeds these parameters to MATCH. They also account for foreground reddening by fitting an additional parameter independently for each spatial cell.
Each cell was fit using four different sets of isochrones: Padova (Marigo et al., 2008; Girardi et al., 2010), BaSTI (Pietrinferni et al., 2004; Cassisi et al., 2005; Pietrinferni, Adriano et al., 2013), PARSEC (Bressan et al., 2012), and MIST (Choi et al., 2016). This resulted in four maps of the SAD (or star formation history) in the PHAT footprint, one for each set of isochrones, each with 16 temporal bins spanning from the present day to the age of the Universe (see Williams et al. 2017 for details). The metallicities used to generate the isochrones were determined not by fitting a single value to each spatial cell, but by modelling the enrichment history of the entire galaxy. Upon finding evidence for chemical enrichment at early times (Mollá et al., 1997), the authors make three radial divisions to M31 and fit an exponentially decreasing (in time) enrichment rate to each region. Their best-fit enrichment decay rates are independently consistent with regions of high star formation and the observed metallicity gradient in M31.
@@ 5 Discussion and Conclusions
Taken together, our detections and upper limits are consistent with a decaying rate of nova formation with delay time; though, strictly speaking, we cannot rule out a uniform formation rate. The maximally likely rate of formation efficiency we recover is, perhaps surprisingly, higher in the final time bin than in the 2-3 Gyr bin, possibly implying a departure from monotonic decay of nova formation. In fact, there is no theoretical expectation that the DTD of novae should be exactly monotonic - the theoretical DTD calculated in Kemp et al. (2021) exhibits several local maxima. However, given the wide range of rates falling within our 1 uncertainty intervals, we caution against over-interpretation of this result.
A decaying formation rate is borne out by simple arguments in stellar evolution. Technically, a binary system can produce novae as soon as the more massive partner evolves into a white dwarf. As delay time increases, the zero-age main sequence masses of both the donor and the degenerate accretor decrease on average, leading to a decrease in formation efficiency of novae (Ritter et al., 1991; Kolb, 1995; Chen et al., 2016) and an increase in recurrence time (Yaron et al., 2005), as less mass is available to transfer to the accretor. Although the PHAT SAD map lacks sufficient resolution at short delay times to probe the “turn-on” phase – the prediction that the formation efficiency of novae should be zero before the formation of the first WDs at delays of 40 Myr –, it does probe, and successfully recover, the broad expectation of a decrease in formation efficiency at longer delays.
Our rough completeness corrections allow us to estimate nova rates (per unit stellar mass) as well as nova counts; these rates are consistent with the shape and height of the Kemp et al. DTD (see Figure 6), with the effective survey length correction yielding a slightly better match than the time-limited sample. To be clear, no adjustments have been made to our data to bring them into alignment with the theoretical rates. This agreement is simply a product of the DTD recovery process outlined in Section 3 and the completeness corrections presented in Section 2.1. In the same figure, we also make a comparison against the Kemp et al. DTD of helium novae – rare events caused either by the deposition of helium, rather than hydrogen, onto a white dwarf, or the fusion of hydrogen into helium on the surface of the dwarf before a nova detonation. As expected, He nova rates are so low as to be inconsistent with our completeness-corrected DTD; this large discrepancy implies that a minute fraction of novae in M31 are He novae.
@@
We conclude with a brief discussion of the comparison between our nova DTD and the DTD for Type Ia SNe recovered by Maoz et al. (2012), which is interesting as some subclasses of novae have been proposed as potential SN Ia progenitors (Maoz et al., 2014). Our completeness-corrected nova DTD is about four orders of magnitude higher than the SN Ia DTD, and noticeably shallower at long delay times. This implies a strong () upper limit on the fraction of nova-producing systems that go on to explode as SN Ia in star-forming galaxies like M31.
Such a discrepancy presents a problem for the single-degenerate theory of Type Ia progenitors, which posits that a significant fraction of SNe Ia arise from gradual mass transfer between a living star and a remnant – the same kind of system that produces novae. If such a small fraction of these systems end their lives as type Ia supernovae (right panel of Figure 6), there may not be enough single-degenerate systems in star-forming galaxies to explain the observed Ia rate (Maoz et al., 2014).
Our work reinforces the previously demonstrated power of DTDs for constraining the progenitor populations of the products of binary stellar evolution. The continuation of high-cadence surveys of nearby, well-studied galaxies such as M31 would provide better constraints on the nova DTD at all lookback times and reduce our reliance on historical nova catalogs of dubious completeness. Resolved stellar populations of those galaxies provide opportunities for the recovery of DTDs for other astrophysical transients. The upcoming extension of the PHAT SAD map to the southern region of M31 (Chen et al., 2025) promises a near-trivial doubling of the nova sample size; once published, this data will represent an exciting opportunity to improve the resolution of the nova DTD.
SAD maps similar to that of Williams et al. (2017) exist for other galaxies in the Local Group (e.g., Lazzarini et al. 2022 for M33, Mazzi et al. 2021; Harris & Zaritsky 2009 for the LMC, Rubele et al. 2018; Harris & Zaritsky 2004 for the SMC), with matching, high-quality catalogs of astrophysical transients and other products of stellar evolution. This research enables further application of the methods described here, which could be leveraged to constrain BPS models, deepening our understanding of key stages of binary stellar evolution and improving the predictions for the rates of rare events such as SNe Ia and black hole mergers.
@@
We acknowledge fruitful discussions with Dan Maoz, Peter Nugent, and Robin Ciardullo. We are especially grateful to Alex Kemp, who kindly provided us with a digital form of his nova DTD. L. Galbany acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, and 2021-SGR-01270. This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided. Specifically, this work used the HTC cluster, which is supported by NIH award number S10OD028483. We acknowledge financial support from NSF grant AST-2307865, and HST grants HST-GO-16741.002-A, HST-GO-16287.002-A, and HST-GO-17179.002-A. The following Python packages were a significant part of this research: Numpy (Harris et al., 2020), Matplotlib (Hunter, 2007), and Astropy (Astropy Collaboration et al., 2013, 2018, 2022).